OpenStudy (anonymous):

Find the x and y intercepts 5x-7y=-35 Please show all of your work.

OpenStudy (anonymous):

for x intercept, y = 0 for y intercept, x = 0

OpenStudy (anonymous):

5x-7y=-35 7y=5x+35 y=5x/7 +5 y intercept x=0, y=-5 (0,5) x intercept y=0, x=-7 (-7,0)

OpenStudy (anonymous):